# MolFORM: Multi-modal Flow Matching for Structure-Based Drug Design

### v2

@@
Abstract
Structure-based drug design (SBDD) seeks to generate molecules that bind effectively to protein targets by leveraging their 3D structural information. While diffusion-based generative models have become the predominant approach for SBDD, alternative non-autoregressive frameworks remain relatively underexplored. In this work, we introduce MolFORM, a novel generative framework that jointly models discrete (atom types) and continuous (3D coordinates) molecular modalities using multi-flow matching. To further enhance generation quality, we incorporate a preference-guided fine-tuning stage based on Direct Preference Optimization (DPO), using Vina score as a reward signal. We propose a multi-modal flow DPO co-modeling strategy that simultaneously aligns discrete and continuous modalities, leading to consistent improvements across multiple evaluation metrics. The code is provided at: https://github.com/huang3170/MolForm.
Machine Learning, ICML
1 Introduction
@@
In recent years, flow matching (Liu et al., 2022b; Lipman et al., 2022) has emerged as a widely studied generative modeling framework. Although flow matching is theoretically equivalent to diffusion models under certain conditions (Gao et al., 2024), empirical performance can vary significantly depending on the choice of scheduling strategy. In image generation tasks, flow matching has been successfully scaled to large datasets and has demonstrated strong performance (Esser et al., 2024; Liu et al., 2023). In the domain of AI for science, especially for molecular and protein generation tasks, researchers have also begun to explore the applicability of flow matching (Jing et al., 2024; Geffner et al., 2025; Campbell et al., 2024). Conceptually, flow matching offers a transport mapping interpretation from the perspective of ordinary differential equations (ODEs), providing a flexible modeling framework that enables task-specific adaptations. For structure-based drug design (SBDD), the generative task involves predicting both atom types and their 3D positions, which can be viewed as a combination of discrete and continuous modalities. Motivated by recent advances, we propose MolFORM, a novel framework for Molecular multi-modal Flow-Optimized Representation Matching. Our results show that carefully designed flow-based models can achieve performance comparable to SOTA diffusion-based approaches.
Furthermore, we incorporate a preference-guided fine-tuning stage using Direct Preference Optimization (DPO), leveraging the Vina score as a chemically informed reward signal to refine the base model. We first propose a discrete flow matching variant of DPO under a uniform noising scheme, which enables preference alignment in the atom type space. More importantly, we demonstrate that the full benefit arises from a multi-flow DPO co-modeling strategy that jointly aligns preferences over both discrete atom types and continuous 3D positions. This additional step leads to notable improvements in the quality of generated molecules, with our experiments showing that the gains from DPO fine-tuning surpass those of standard diffusion-based models. To the best of our knowledge, this is the first work to apply preference alignment in a multi-flow setting, highlighting its potential for broader adoption in similar generative tasks.
2 Related work
2.1 Structure-Based Drug Design.
@@ 2 Related work
With the increasing availability of structural data, generative models have attracted significant attention for structure-based molecule generation. Early methods (Skalic et al., 2019) utilized sequence generative models to produce SMILES representations from protein contexts. Driven by advancements in 3D geometric modeling, subsequent studies directly generate molecules in 3D space. For instance, Ragoza et al. (2022) employ voxelized atomic density grids within a Variational Autoencoder framework. Other methods use autoregressive models to sequentially place atoms or chemical groups (Luo et al., 2021; Peng et al., 2022), while FLAG (Zhang et al., 2023) and DrugGPS (Zhang & Liu, 2023) leverage chemical priors to construct realistic ligand fragments incrementally. More recently, diffusion models have demonstrated notable success by progressively denoising atom types and coordinates, maintaining SE(3)-equivariant symmetries (Guan et al., 2023; Lin et al., 2022; Schneuing et al., 2023; Huang et al., 2023; Guan et al., 2024; Zhang et al., 2024). Despite these advances, existing models often struggle with generating molecules simultaneously optimized for multiple desirable properties, such as binding affinity, synthesizability, and low toxicity—critical considerations in drug discovery (D Segall, 2012).
2.2 Flow Matching.
@@
Flow matching (Lipman et al., 2022; Liu et al., 2022b) is a continuous-time generative modeling framework that generalizes diffusion models by learning a vector field to transport a simple prior $q(\bm{x})$ toward the data distribution $p_{\text{data}}(\bm{x})$. It defines a conditional path $p_{t}(\bm{x}\mid\bm{x}_{1})$ that interpolates between $q(\bm{x})$ and the target $\delta(\bm{x}-\bm{x}_{1})$, and learns the marginal vector field $v(\bm{x},t)=\mathbb{E}_{\bm{x}_{1}\sim p_{t}(\bm{x}_{1}\mid\bm{x})}[u_{t}(\bm{x}\mid\bm{x}_{1})]$ using a neural network ${v}_{\theta}(\bm{x},t)$. Diffusion models can be seen as a special case under Gaussian interpolation (Gao et al., 2024). Flow matching has demonstrated strong performance in image and video generation (Liu et al., 2023; Esser et al., 2024), and is gaining traction in scientific domains such as protein generation  (Campbell et al., 2024), protein conformation generation (Jing et al., 2024), where it naturally handles both discrete atom types and continuous 3D coordinates.
2.3 Preference alignment of generative models.
@@
While maximizing data likelihood is standard in generative modeling, it often fails to align with downstream user preferences. Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2020; Ouyang et al., 2022) and its variants have been widely adopted to align large language models with human intent. Recent efforts extend these ideas to diffusion models, treating generation as a multi-step decision process (Uehara et al., 2024; Wallace et al., 2023). Direct Preference Optimization (DPO) (Rafailov et al., 2023) offers a simpler alternative by bypassing reinforcement learning and directly optimizing models against pairwise preference data. DPO has shown competitive results in both language and image domains (Wallace et al., 2023; Zhou et al., 2024). Meanwhile, we also observe that some studies (Gu et al., 2024; Cheng et al., 2024) have begun to incorporate preference alignment into structure-based drug design (SBDD), highlighting its potential to improve biological plausibility and design success. However, applying DPO to jointly model discrete and continuous variables remains largely unexplored.
Figure 1: Overview of MolFORM. This workflow can be summarized as two steps: 1) Employs multi-flow generation to construct the base model. 2) Applies DPO to fine-tune the dual modalities, using the Vina score as the reward.
3 Methods
3.1 Problem definitions.
@@ 3 Methods
3.2 Multi-modal Flow matching
@@
The Conditional Flow Matching (CFM) framework (Lipman et al., 2022; Liu et al., 2022b) learns a time-dependent flow $\psi_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that transports samples from a source distribution $p_{0}$ to a target distribution $p_{1}$, governed by the ODE $\frac{d}{dt}\mathbf{x}_{t}=u_{t}(\mathbf{x}_{t})$ with $\mathbf{x}_{t}=\psi_{t}(\mathbf{x}_{0})$. Since the exact vector field $u_{t}$ is generally intractable, the conditional formulation enables a tractable objective by conditioning on both source and target samples, where the $v_{\theta}$ is a time- depended, learnable model with parameters $\theta$:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\left\|v_{\theta}(\mathbf{x}_{t},t)-u_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{1})\right\|^{2}.$ (1)
In our structure-based drug design (SBDD) task, we model each atom type as a categorical variable, where $x\in\mathcal{K}$ and $\mathcal{K}=\{1,\dots,K\}$ is the set of possible atom types with cardinality $K$. To model the generative process over atom types, we adopt the Discrete Flow Matching (DFM) framework (Campbell et al., 2024), which constructs a time-dependent probability flow $p_{t}(v)$ that evolves from a known noise distribution $p_{0}(v)$ to the empirical data distribution $p_{1}(v)=p_{\mathrm{data}}(v)$ over a time interval $t\in[0,1]$.
This flow is defined via a family of conditional distributions $\pi_{t}(v_{t}\mid v_{1})$, referred to as conditional flows, where $v_{1}\in[k]$ is the true atom type and $v_{t}$ is its corrupted version at time $t$. We adopt the uniform corruption process:
$\pi_{t}(v_{t}\mid v_{1})=(1-t)\cdot\mathrm{Uniform}([k])+t\cdot\delta_{v_{1}}(v_{t}),$ (2)
where $\delta_{v_{1}}(v_{t})$ is the Kronecker delta function (equal to 1 if $v_{t}=v_{1}$, and 0 otherwise), and $\mathrm{Uniform}([k])=1/k$ for all $v_{t}\in[k]$. This defines a simple interpolation between a uniform noise distribution $p_{0}(v)=\mathrm{Uniform}([k])$ at $t=0$ and the data distribution $p_{1}(v)=p_{\mathrm{data}}(v)$ at $t=1$. The marginal distribution at time $t$ is given by averaging over the data distribution:
$p_{t}(v_{t})=\mathbb{E}_{v_{1}\sim p_{\mathrm{data}}}\left[\pi_{t}(v_{t}\mid v_{1})\right].$ (3)
To train the model, we learn a denoising distribution $p^{\theta}(v^{i}_{1}\mid\mathbf{v}_{t})$ that predicts the original atom type $v^{i}_{1}\in[k]$ from a corrupted ligand type sequence $\mathbf{v}_{t}=\{v^{i}_{t}\}_{i=1}^{N_{M}}$. The model is optimized using a standard cross-entropy loss:
@@
$\mathcal{L}_{\mathrm{CE}}=\mathbb{E}_{t,\,\mathbf{v}_{1},\,\mathbf{v}_{t}\sim\pi_{t}(\cdot\mid\mathbf{v}_{1})}\left[-\sum_{i=1}^{N_{M}}\log p^{\theta}(v^{i}_{1}\mid\mathbf{v}_{t})\right],$ (4)
where $\pi_{t}(v^{i}_{t}\mid v^{i}_{1})$ denotes the uniform corruption process.
To improve training stability, especially when $t\to 1$, we adopt a reparameterized formulation inspired by diffusion-based objectives. Instead of regressing the velocity field directly, we predict the target sample $\hat{\mathbf{x}}_{1}=v_{\theta}(\mathbf{x}_{t},t)$ and define the loss in terms of reconstructed flows:
$\mathcal{L}_{\text{reparam}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\left\|u_{t}(\mathbf{x}_{t}|\hat{\mathbf{x}}_{1},\mathbf{x}_{0})-u_{t}(\mathbf{x}_{t}|\mathbf{x}_{1},\mathbf{x}_{0})\right\|^{2}.$ (5)
On Euclidean manifolds, where $u_{t}(\mathbf{x}_{t}|\mathbf{x}_{1},\mathbf{x}_{0})=\mathbf{x}_{1}-\mathbf{x}_{0}$, this reduces to a simple mean squared error:
$\mathcal{L}_{\text{pos}}=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{1}}\left\|\hat{\mathbf{x}}_{1}-\mathbf{x}_{1}\right\|_{2}^{2}.$ (6)
Similarly, we reparameterized the cross-entropy loss to predict the clean sample $\hat{\mathbf{v}}_{1}$ and match the conditional flows induced by $\hat{\mathbf{v}}_{1}$ and $\mathbf{v}_{1}$. Here the loss function is expressed as:
$\mathcal{L}_{\text{type}}=\mathbb{E}_{t,\,\mathbf{v}_{0},\mathbf{v}_{1}}\left[\mathrm{CE}\left(\hat{\mathbf{v}}_{1},\,\mathbf{v}_{1}\right)\right],$ (7)
This reparameterization has been shown to enhance numerical stability, especially when sampling with large $t$.
Figure 2: Visualizations of reference molecules and generated ligands for protein pockets (4yhj) generated by Reference, Targetdiff, MolFORM and MolFORM-DPO. Vina score, QED, and SA are reported below.
Chamfer loss

### v3

# MolFORM: Preference-Aligned Multimodal Flow Matching for Structure-Based Drug Design

@@
Abstract.
Structure-based drug design (SBDD) aims to efficiently discover high-affinity ligands within vast chemical spaces. However, current generative models struggle with objective misalignment and rigid sampling budgets. We present MolFORM, a fast multi-modal flow matching framework for discrete atom types and continuous coordinates. Crucially, to bridge the gap between generative capability and biochemical objectives, we introduce two distinct post-training strategies: (1) Direct Preference Optimization (DPO), which performs offline alignment using ranked preference pairs; and (2) an online reinforcement learning paradigm that optimizes the generative flow directly on the forward process. Both strategies effectively navigate the chemical space toward high-affinity regions. MolFORM achieves state-of-the-art results on CrossDocked2020 benchmark (Vina Score -7.60, Diversity 0.75), demonstrating that incorporating preference alignment mechanisms—whether via offline optimization or online reinforcement—is crucial for steering generative models toward high-affinity binding regions. The source code for MolFORM is publicly available at https://github.com/daiheng-zhang/SBDD-MolFORM.
Structure based drug design, Flow matching, Preference alignment
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
In recent years, flow matching (Liu et al., 2022b; Lipman et al., 2022) has emerged as a widely studied generative modeling framework. Although flow matching is theoretically equivalent to diffusion models under certain conditions (Gao et al., 2024), empirical performance can vary significantly depending on the choice of scheduling strategy. In image generation tasks, flow matching has been successfully scaled to large datasets and has demonstrated strong performance (Esser et al., 2024; Liu et al., 2023). In the domain of AI for science, especially for molecular and protein generation tasks, researchers have also begun to explore the applicability of flow matching (Jing et al., 2024; Geffner et al., 2025; Campbell et al., 2024). Conceptually, flow matching offers a transport mapping interpretation from the perspective of ordinary differential equations (ODEs), providing a flexible modeling framework that enables task-specific adaptations. For structure-based drug design (SBDD), the generative task involves predicting both atom types and their 3D positions, which can be viewed as a combination of discrete and continuous modalities. Motivated by recent advances, we propose MolFORM, a novel framework for Molecular multi-modal Flow-Optimized Representation Matching. To further enhance sampling efficiency, we also designed an auxiliary confidence head capable of predicting confidence scores for generated structures, serving as a basis for ranking high-quality candidates.
Furthermore, to bridge the gap between generative capability and biochemical objectives, we introduce the preference-guided fine-tuning stage comprising two distinct strategies: offline Direct Preference Optimization (DPO) (Rafailov et al., 2023; Wallace et al., 2023) and online Reinforcement Learning (RL) (Zheng et al., 2025). We demonstrate that the substantial performance gains stem from a multi-flow co-modeling strategy that jointly aligns preferences over both discrete atom identities and continuous 3D positions. Beyond offline alignment, we further incorporate an online RL paradigm that optimizes the generative flow directly on the forward process. By dynamically contrasting positive and negative generations sampled during training, this approach efficiently navigates the chemical space toward high-affinity regions. Our experiments show that these rigorous flow alignment techniques leverage the Vina score (Trott and Olson, 2010) as a chemically informed reward signal to significantly enhance molecule quality, while enabling the flexible design of reward functions tailored to specific requirements for the target molecule.
2. Related work
2.1. Structure-Based Drug Design.
@@ 2. Related work
With the increasing availability of structural data, generative models have attracted significant attention for structure-based molecule generation. Early methods (Skalic et al., 2019) utilized sequence generative models to produce SMILES representations from protein contexts. Driven by advancements in 3D geometric modeling, subsequent studies directly generate molecules in 3D space. For instance, Ragoza et al. (2022) employ voxelized atomic density grids within a Variational Autoencoder framework. Other methods use autoregressive models to sequentially place atoms or chemical groups (Luo et al., 2021; Peng et al., 2022), while FLAG (Zhang et al., 2023) and DrugGPS (Zhang and Liu, 2023) leverage chemical priors to construct realistic ligand fragments incrementally. More recently, diffusion models have demonstrated notable success by progressively denoising atom types and coordinates, maintaining SE(3)-equivariant symmetries (Guan et al., 2023; Lin et al., 2022; Schneuing et al., 2023; Huang et al., 2023; Guan et al., 2024; Zhang et al., 2024). Despite these advances, existing models often struggle with generating molecules simultaneously optimized for multiple desirable properties, such as binding affinity, synthesizability, and low toxicity in drug discovery (D Segall, 2012).
2.2. Flow Matching.
@@
Flow matching (Lipman et al., 2022; Liu et al., 2022b) is a continuous-time generative modeling framework that generalizes diffusion models by learning a time-dependent vector field to transport a simple prior $q(\bm{x})$ toward the data distribution $p_{\text{data}}(\bm{x})$. It defines a conditional probability path $p_{t}(\bm{x}\mid\bm{x}_{1})$ that interpolates between $q(\bm{x})$ and the target $\delta(\bm{x}-\bm{x}_{1})$, and learns the corresponding marginal vector field $v(\bm{x},t)=\mathbb{E}_{\bm{x}_{1}\sim p_{t}(\bm{x}_{1}\mid\bm{x})}[u_{t}(\bm{x}\mid\bm{x}_{1})]$ using a neural network $v_{\theta}(\bm{x},t)$. Diffusion models can be viewed as a special case of flow matching under Gaussian interpolation (Gao et al., 2024). As a simulation-free training paradigm for continuous normalizing flows, flow matching offers flexibility in choosing probability paths and time schedules, while recent method such as Rectified Flow encourage straighter transport trajectories and enable efficient sampling with fewer ODE steps. Flow matching has demonstrated strong performance in large-scale image and video generation (Liu et al., 2023; Esser et al., 2024), and is increasingly adopted in scientific domains such as protein generation (Campbell et al., 2024) and protein conformation modeling (Jing et al., 2024). Moreover, discrete flow models extend flow matching to discrete state spaces via continuous-time Markov chains, recovering discrete diffusion as a special case and enabling unified modeling of multimodal settings that couple discrete atom types with continuous 3D structures.
2.3. Preference Alignment of Diffusion Models.
While maximizing data likelihood is standard in generative modeling, it often fails to align with downstream user preferences. Reinforcement learning from human feedback (RLHF) (Ziegler et al., 2020; Ouyang et al., 2022) has been widely adopted to align large language models with human intent. Recent efforts extend these ideas to diffusion models by treating generation as a multi-step decision process (Uehara et al., 2024; Wallace et al., 2023). To implement this, early works typically formulate sampling as a Markov Decision Process (MDP), discretizing the reverse process to apply Policy Gradient algorithms (Black et al., 2023; Fan et al., 2023). Notably, methods like FlowGRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025) successfully adapt Group Relative Policy Optimization (GRPO) to diffusion by leveraging SDE-based stochasticity for exploration. However, these reverse-process approaches often suffer from solver restrictions and forward-reverse inconsistency. To address these limitations, DiffusionNFT (Zheng et al., 2025) introduces a novel online RL paradigm that optimizes directly on the forward process via flow matching. By contrasting positive and negative generations to define an implicit improvement direction, DiffusionNFT eliminates the need for likelihood estimation and achieves significantly higher training efficiency compared to GRPO-based methods.
@@
Alternatively, Direct Preference Optimization (DPO) (Rafailov et al., 2023) offers a simpler paradigm by bypassing reinforcement learning and directly optimizing models against pairwise preference data. DPO has shown competitive results in both language and image domains (Wallace et al., 2023; Zhou et al., 2024b). In the context of structure-based drug design (SBDD), recent studies (Gu et al., 2024; Cheng et al., 2024) have also begun to incorporate preference alignment to improve biological plausibility and design success.
3. Methods
3.1. Problem definitions.
@@ 3. Methods
Figure 1. Overview of MolFORM. This workflow can be summarized as two steps: 1) Employs multi-flow generation to construct the base model. 2) Applies DPO to fine-tune the dual modalities, using the Vina score as the reward.
3.2. Multi-modal Flow Matching
The overall framework of MolFORM is illustrated in Figure 1. We model a ligand as a multimodal object consisting of continuous atomic coordinates $x\in\mathbb{R}^{N_{M}\times 3}$ and discrete atom types $v\in[K]^{N_{M}}$, conditioned on a protein pocket $p$. MolFORM jointly learns (i) a continuous flow for coordinates and (ii) a discrete flow for atom types, using a shared SE(3)-equivariant backbone and synchronized time $t\in[0,1]$.
Continuous flow matching for atomic coordinates.
@@ 3.2. Multi-modal Flow Matching
Conditional Flow Matching (CFM) learns a time-dependent flow $\psi_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that transports samples from a source distribution $p_{0}$ to a target distribution $p_{1}$, governed by the ODE $\frac{d}{dt}x_{t}=u_{t}(x_{t})$ with $x_{t}=\psi_{t}(x_{0})$. Since the exact marginal vector field $u_{t}$ is generally intractable, CFM defines a conditional vector field $u_{t}(x_{t}\mid x_{0},x_{1})$ and trains a neural vector field (velocity) $\mathbf{v}_{\theta}$ by regression:
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,\,x_{0},\,x_{1}}\left\|\mathbf{v}_{\theta}(x_{t},v_{t},p,t)-u_{t}(x_{t}\mid x_{0},x_{1})\right\|_{2}^{2},$
where we omit conditioning variables in $u_{t}(\cdot)$ for brevity.
In this work, we use the rectified flow on Euclidean space:
$x_{t}=(1-t)x_{0}+tx_{1},\qquad u_{t}(x_{t}\mid x_{0},x_{1})=x_{1}-x_{0},$
with $x_{0}\sim p_{0}$ a Gaussian prior and $x_{1}\sim p_{\mathrm{data}}$ a data sample.
Discrete flow matching for atom types.
For discrete atom types, we adopt the Discrete Flow Matching based on continuous-time Markov chains (CTMCs). We define a family of conditional flows $\pi_{t}(v_{t}\mid v_{1})$ and use uniform corruption:
$\pi_{t}(v_{t}\mid v_{1})=(1-t)\cdot\mathrm{Uniform}([K])+t\cdot\delta_{v_{1}}(v_{t}),$
where $\delta_{v_{1}}(v_{t})$ is the Kronecker delta and $\mathrm{Uniform}([K])=1/K$. The corresponding marginal at time $t$ is
$p_{t}(v_{t})=\mathbb{E}_{v_{1}\sim p_{\mathrm{data}}}\big[\pi_{t}(v_{t}\mid v_{1})\big].$
We train a denoising posterior $p_{\theta,1|t}(v_{1}\mid v_{t},p,t)$ to predict clean atom types. A standard cross-entropy objective is
$\mathcal{L}_{\mathrm{CE}}=\mathbb{E}_{t,\,v_{1},\,v_{t}\sim\pi_{t}(\cdot\mid v_{1})}\left[-\sum_{i=1}^{N_{M}}\log p_{\theta,1|t}\!\left(v^{i}_{1}\mid v_{t},p,t\right)\right],$
where $v_{1}=\{v_{1}^{i}\}_{i=1}^{N_{M}}$ and $v_{t}=\{v_{t}^{i}\}_{i=1}^{N_{M}}$.
Reparameterized training with $x$-prediction.
Although $\mathbf{v}_{\theta}$ is the velocity field used for sampling, we adopt a reparameterized training objective that computes regression in $x$-space for improved numerical stability. Given the current noisy state $(x_{t},v_{t})$, we form a one-step endpoint reconstruction from time $t$ to $1$:
$\hat{x}_{1}\;=\;x_{t}+(1-t)\,\mathbf{v}_{\theta}(x_{t},v_{t},p,t).$
We then define the reconstructed-flow loss
$\mathcal{L}_{\mathrm{reparam}}(\theta)=\mathbb{E}_{t,\,x_{0},\,x_{1}}\left\|u_{t}(x_{t}\mid\hat{x}_{1},x_{0})-u_{t}(x_{t}\mid x_{1},x_{0})\right\|_{2}^{2}.$
On Euclidean manifolds with the straight-line path $u_{t}(x_{t}\mid x_{0},x_{1})=x_{1}-x_{0}$, this reduces to a simple $x$-space MSE:
$\mathcal{L}_{\mathrm{pos}}=\mathbb{E}_{t,\,x_{0},\,x_{1}}\left\|\hat{x}_{1}-x_{1}\right\|_{2}^{2}.$
Similarly, for discrete atom types we directly predict the clean posterior $p_{\theta,1|t}(\cdot\mid v_{t},p,t)$ and use the corresponding cross-entropy (this fixes the $v_{0}$$v_{t}$ notation issue):
@@
$\mathcal{L}_{\mathrm{type}}=\mathbb{E}_{t,\,v_{1},\,v_{t}\sim\pi_{t}(\cdot\mid v_{1})}\left[\mathrm{CE}\!\left(p_{\theta,1|t}(\cdot\mid v_{t},p,t),\,v_{1}\right)\right].$
Chamfer loss.